zarasht:

The reaction A + 2B + C → Products has a rate law of Rate = k[A]². By what factor would the reaction rate change if is tripled?

justjm:

Same logic as a question I believe I solved before \(R=k\left[A\right]^{2}\left[B\right]\) The concentration of B is now tripled, so you can plug in 3B \(R=k\left[A\right]^{2}\left[3B\right]\) And you can simplify to find the new R. Let me know if you have any questions.